let z be the set of integers and R be the relation on Z defined as: aRb if and only if 1+ab>0 then


1
Expert's answer
2021-08-03T13:16:50-0400

Given relation is "aRb" is "1+ab>0."

Considering both "a" and "b" are real numbers, we know that "ab=ba"


"aRb=1+ab>0=>bRa=1+ba=1+ab>0"

Then "R" is a symmetric relation.



"aRa=1+a^2>0"

Then "R" is a reflexive relation.


Let "a=0.5, b=-0.5," and "c=-4." Then


"aRb=1+ab=1+0.5(-0.5)=0.75>0"

"bRc=1+bc=1+(-0.5)(-4)=3>0"

But


"aRc=1+ac=1+0.5(-4)=-1<0"

"aRc" is not a relation.


Hence "R" is not a equivalence relation, but is a reflexive and symmetric relation.


reflexive and symmetric relation.